OpenStudy (anonymous):

if x has a poisson distribution with u=3, what is the probability that x=3

OpenStudy (anonymous):

\[P(x=3)=\frac{3^3e^{-3}}{3!}\]if i remember correctly

OpenStudy (anonymous):

ok that was wrong general form is \[P(x=k)=\frac{\lambda^ke^{-\lambda}}{k!}\]

OpenStudy (anonymous):

compute via a calculator, although \[\frac{3^3}{3!}=\frac{3^2}{2}=\frac{9}{2}\] you will still need a calculator for \(e^{-3}\) though